How many grams of water are contained in 2.78g og CaSo4*2H2O
1
Expert's answer
2012-10-08T10:22:48-0400
Molecular mass of CaSO4*2H2O is 172 g/mol. The one mole of CaSO4*2H2O contains
two moles of water. The amount of CaSO4*2H2O is
n(CaSO4*2H2O)=m(CaSO4*2H2O)/M(CaSO4*2H2O)=2.78 g / 172 g/mol = 0.0162 mol.
n(H2O)=2*n(CaSO4*2H2O)=0.0323 mol. m(H2O)=n(H2O)*M(H2O)=0.0323*18=0.582 g.